OpenStudy (anonymous):

Anyone good with log properties? If ln(3x+5) + ln(3x-5) = 6 + ln4 I can use prop 4 to get ln(3x+5)(3x-5)=6+ln4 Does (3x+5)(3x-5)=10?

myininaya (myininaya):

i would write \[\ln(3x+5)+\ln(3x-5)-\ln4=6\] and use properties to write ln(a)+ln(b)-ln(c) as ln(a*b/c)

myininaya (myininaya):

and I would write \[6 \] as \[\ln(e^6)\]

myininaya (myininaya):

so then you would have \[\frac{a*b}{c}=e^6\] see if you can figure it out from here let me know if you have any questions

OpenStudy (anonymous):

I set it up like that, but I still don't understand what happens next.... what do I do with the e?

myininaya (myininaya):

what do you mean its just a number like 5 or 3.4422 or 2 you want to solve for x

myininaya (myininaya):

\[\frac{(3x+5)(3x-5)}{4}=e^6\]

myininaya (myininaya):

\[(3x+5)(3x-5)=4e^6\] \[(3x+5)(3x-5)=4e^6\] \[9x^2-25=4e^6\] \[9x^2=4e^6+25\] you try the rest